OpenStudy (cutiecomittee123):

A triangle has one sisde with length 9 and another side with length 6. The angle opposite the side of length 6 measures 40 degrees. What is the measure of the angle opposite the side length 9? I know we have to use law of sines. Also in the answer you should incude how many solutions are possible. In this case I think it is 2 answers since b/asinA<1. Therefore I am pretty sure we use b1-sin^-1(b/asinA) and b2=180 degrees-sin^-1(b/asinA)

OpenStudy (owlcoffee):

Well, you have to set up the law of sines: \[\frac{ a }{ \sin \angle A }=\frac{ b }{ \sin \angle B }\] \[\frac{ 6 }{ \sin 40 }=\frac{ 9 }{ \sin \angle B }\]

OpenStudy (owlcoffee):

Remember also: \[\sin \alpha = \sin(180-\alpha)\]

OpenStudy (cutiecomittee123):

I have all of that so far,plus I know SinB=9/6Sin40

OpenStudy (anonymous):

final answer is what you get with a calculator

OpenStudy (anonymous):

take the arcsine of your result

OpenStudy (anonymous):

sup cutie xd

OpenStudy (cutiecomittee123):

@satellite73 well I plugged it into a calulator and I got 1.12 then i took the inverse sine but my calculator says error

OpenStudy (cutiecomittee123):

nevermind I retried it and it came out with .96, I took the inverse and I got SinB is approximately 73.74